# Viscoplasticity can stabilise liquid collar motion on vertical cylinders

### v2

@@ IV.3 Intermediate regions
$h=B^{1/2}\frac{V}{4\pi}\hat{X}^{2}+B\left(\hat{A}_{\pm}-\frac{V\hat{X}^{4}}{48\pi}\right)+B^{5/4}\hat{h}_{\pm}(\hat{X})+\dots,\quad Y=B^{1/2}\hat{Y}_{\pm}(\hat{X})+\dots,$ (53)
@@
$-6\mathcal{U}V\hat{X}^{2}\left(\frac{V\hat{X}^{2}}{4\pi}-\hat{Y}_{\pm}\right)=\mathcal{J}\hat{Y}^{2}_{\pm}(3V\hat{X}^{2}-4\pi\hat{Y}_{\pm}).$ (55)
The cubic equation (55) can be solved to find $\hat{Y}_{\pm}$ and (54) can then be integrated to find $\hat{h}_{\pm}$. In the far field approaching the main collar body, we expect $Y\ll h$, so $\hat{Y}_{\pm}\ll\hat{X}^{2}$ as $\hat{X}\rightarrow\mp\infty$. Therefore, (55) implies
@@ IV.4 Bretherton regions
$h=B\mathcal{H}_{\pm}(\xi)+\dots,\quad Y=B\mathcal{H}_{\pm}(\xi)+\dots,$ ($)61𝑎𝑏$
where the subscripts again distinguish quantities in the rear and front Bretherton regions. The evolution equation (30) again says that the leading-order flux is constant, with $q=-B^{5/2}\mathcal{U}H_{\infty}^{+}$ in the rear Bretherton region, and $q=-B^{5/2}\mathcal{U}H_{\infty}^{-}$ in the front. Therefore, (30) and (($)61𝑎𝑏$) give
@@
We introduce rescaled variables, defining $X_{\pm}=(3|\mathcal{U}|)^{1/3}\xi/{H_{\infty}^{\pm}}$, and let ±$\mathcal{H}_{\pm}=H_{\infty}^{\pm}H_{\pm}(X_{\pm})$, so that (62) becomes
@@
$h\sim B^{1/2}H_{2}\hat{X}^{2}\frac{\left(3|\mathcal{U}|\right)^{2/3}}{2H_{\infty}^{+}}+BH_{\infty}^{+}H_{0}+B^{5/4}\frac{2(H_{\infty}^{+})^{2}}{3H_{2}^{2}(3|\mathcal{U}|)^{1/3}\hat{X}}+\dots$ (66)
@@
$h\sim B^{1/2}\frac{V}{4\pi}\hat{X}^{2}+B\hat{A}_{+}-B^{5/4}\frac{8\pi^{2}\mathcal{U}}{3V^{2}\hat{X}}+\dots.$ (67)
@@
$h\sim B^{1/2}\frac{V}{4\pi}\hat{X}^{2}+B\left(\hat{A}_{+}-2\pi+\frac{4\pi^{2}\mathcal{J}}{V}\right)-B^{5/4}\frac{8\pi^{2}\mathcal{U}}{3V^{2}\hat{X}}+\dots.$ (70)
@@
We have now derived a leading-order expression for $\mathcal{U}$, and relations between $H_{\infty}^{\pm}$, $\mathcal{U}$, $\mathcal{J}$ and $V$. There are further details of the asymptotic model, discussed in appendices B and C, which are required to expose the full structure of the solution. In the front Bretherton region, the layer height, $\mathcal{H}$, has an undulating shape, and there is a sequence of points within the Bretherton region around which ′′′$\mathcal{H}^{\prime\prime\prime}\approx 0$. An asymptotically short inner region, in which $Y$ is no longer equal to $h$ at leading order, must exist around each point where ′′′$\mathcal{H}^{\prime\prime\prime}=0$. We discuss the structure of these inner regions in appendix B. At the outer edges of both Bretherton regions, $Y$ must approach zero in order for the solution to match onto the uniform unyielded film away from the collar. In appendix C, we discuss the structure of two ‘rigidification regions’ near to the points $\zeta=\zeta_{0}$ and $\zeta=\zeta_{1}$, where $Y\rightarrow 0$. We find that in the rigidification regions, $\zeta-\zeta_{1}=O(B^{1/2})$ and $\zeta_{0}-\zeta=O(B^{1/2})$, respectively, but with $\zeta-\zeta_{1}\ll B^{1/2}\log(1/B)$ and $\zeta_{0}-\zeta\ll B^{1/2}\log(1/B)$, and that the Bretherton region solutions outlined above strictly hold when $\zeta-\zeta_{1}\gg B^{1/2}\log(1/B)$ and $\zeta_{0}-\zeta\gg B^{1/2}\log(1/B)$, respectively. This motivates the $O(B^{1/2}\log(1/B))$ length scales identified in figure 3(b,c) for the distance between the Bretherton regions and the points $\zeta=\zeta_{0}$ and $\zeta=\zeta_{1}$. None of these details impact the leading-order predictions for $\mathcal{U}$ and $H_{\infty}^{+}$ determined above.
@@ V Translating collars at small $B$
The asymptotic model outlined in §IV provides information on how a single, isolated collar propagates into a uniform film ahead of it when $B\ll 1$. Numerical solutions of (22), such as the one shown in figure 2, can also describe the motion of a translating collar when they are run to sufficiently long times to allow for a collar to develop and then begin steadily propagating. Since we enforce periodicity at the sides of our computational domain when solving (22) numerically, the film thicknesses at the front edge and rear edge of a translating collar must be the same. This means that the numerical simulations always reach a state at late times in which the collar is steadily translating. Therefore, we can directly compare predictions from the asymptotic model, in the case that $V$ and $\mathcal{U}$ are constant and $H_{\infty}^{+}=H_{\infty}^{-}$, against late-time predictions from numerical simulations: we do so in §V.1. In §V.2, we discuss how an isolated collar evolves in an unsteady manner when the film thickness ahead and behind are unequal, which can be described by the asymptotic model.
@@ V.1 Steady translation
Figure 4: Dependence of collar volume and speed on Bingham number, $\mathcal{J}$, for fixed values of $H_{\infty}^{-}$. Variation of (a) $V$, and (b) $\mathcal{U}$, for $10<\mathcal{J}<$ 1, according to (77) and (78). Each solid line corresponds to one value of $H_{\infty}^{-}$, with the values shown being $H_{\infty}^{-}=\{0.4,0.7,1,1.3,1.6\}$. The large arrows indicate increasing values of $H_{\infty}^{-}$. The dashed line in (a) is $V=2\pi\mathcal{J}$, which corresponds to the minimum volume required for the collar to translate.
@@
Figure 5: Late-time, steady-state collar speed, $\mathcal{U}=B^{-3/2}U$, from numerical solutions of (22), with initial condition (29) and domain length $\mathcal{L}=2\pi/k_{\mathrm{m}}$, for various values of the Bond number, $B$, and the Bingham number, $\mathcal{J}=J/B$. Simulations are run to $t=3000$, and $U$ is approximated by fitting a straight line to the location of $\max_{z}H$ during the interval $2800\leq t\leq 3000$. The solid black line shows the leading-order asymptotic prediction (78) for $\mathcal{U}$, where we approximate the collar volume by the total volume in the domain, i.e., $V=\mathcal{L}$. The shaded region is $\mathcal{J}>1$, where an initially flat layer is stable to small perturbations. The dashed line shows the collar speed determined via (D.18), the global force balance in which the $O(B^{5/4}\log{B})$ correction to the wall shear stress is included, where we take $B=0.025$. The dotted line shows the collar speed determined via (80) and (81), the global force balance in the case that $J=O(B^{3/2})$, where again we take $B=0.025$.
@@
Rather than treating $H_{\infty}^{-}$ as fixed, as we do in figure 4, we could instead treat $V$ as fixed, considering a collar with fixed volume. In this case, from (77) and (78), both the film thickness, $H_{\infty}^{-}$, and the propagation speed, $|\mathcal{U}|$, decrease as $\mathcal{J}$ is increased. The case of fixed $V$ is more amenable to direct comparison with numerical solutions of (22) in periodic domains, since the total volume of fluid within the periodic domain is fixed to be $\mathcal{L}$. The asymptotic prediction for $\mathcal{U}$ compares well with late-time collar speeds in numerical simulations (figure 5, solid line). In figure 5, for $\mathcal{J}<1$, there is good (but not perfect) agreement between $B^{-3/2}U$ from numerical simulations, and $\mathcal{U}$, which provides evidence to support the scalings and solutions in the asymptotic theory for small $B$. Even when $B$ is not very small, for example when $B=0.2$ (figure 5), there is reasonably good quantitative agreement between the numerical and asymptotic predictions for translation speed across $10\leq\mathcal{J}<$ 1, although the asymptotics generally predict a modestly faster speed. At smaller values of $\mathcal{J}$, the speeds from numerical solutions appear to converge towards the asymptotic prediction as $B$ is decreased, but at larger $\mathcal{J}$ there is not clear convergence towards the leading-order asymptotic prediction for the values of $B$ tested in figure 5. When $\mathcal{J}$ is large enough that viscoplastic effects are significant, we may expect that convergence to the small-$B$ asymptotic regime only occurs at much smaller $B$ than when $\mathcal{J}\approx 0$. This is because the largest error term in $Y$ in the main body of the collar is only expected to be $O(B^{1/4})$ larger than the leading-order term (($)39𝑎𝑏$), so $B$ must be very small before the error in $Y$, and by extension the error in the viscous drag in the collar, is negligible. For numerical simulations with $\mathcal{J}>1$, there is generally no motion, or very minimal motion, in the layer, reflected by the data showing $U=0$ for $\mathcal{J}>1$ in figure 5. This provides confirmation of the stability threshold identified in the linear stability analysis, §III.1. However, the asymptotic prediction for $\mathcal{U}$ extends beyond $\mathcal{J}=1$, with $\mathcal{U}\rightarrow 0$ as $\mathcal{J}\rightarrow V/2\pi$, which indicates that it could be possible to observe a translating collar when $1\leq\mathcal{J}<V/2\pi$ if a sufficiently large perturbation were applied to the free surface height initially to trigger instability. Here, we focus on the more natural case in which the layer is flat, or very close to flat, initially, in numerical simulations. The asymptotic solution does not exist for $\mathcal{J}\geq V/2\pi$, since in that case, the weight of the collar is no longer sufficient to make the fluid yield so there can be no translation of the collar.
@@
where the $O(B^{5/4}\log(1/B))$ correction arises from contributions from the wall shear stress in the intermediate regions. Replacing the left-hand side of (76) with the expression (79) yields a prediction for the collar speed which generally agrees more closely with numerical results than the leading-order prediction for $\mathcal{U}$ (figure 5, dashed line).
@@
is a good approximation for $200<\Gamma<$ 20. Numerical evaluation of $\tilde{F}(\Gamma)$ suggests that $\tilde{F}\approx 2\pi$ when $\Gamma$ is sufficiently large, suggesting that the result (76) for $J=O(B)$ can be recovered in the limit $\hat{J}\rightarrow\infty$. The collar speed determined via (80), with the approximation (81), yields an improved estimate of the steady collar speed when $\mathcal{J}\ll 1$ (figure 5, dotted line), compared to the leading-order prediction (76) for $\mathcal{U}$. For $B=0.025$, $\Gamma<20$ holds for $0<\mathcal{J}\leq 0.4$, so use of (81) is justified at least for $0<\mathcal{J}\leq 0.4$ in figure 5.
@@ VI Discussion
Unlike in the problem without gravity [15, 16], an initially flat layer of liquid can be linearly unstable to infinitesimal perturbations in the free-surface height, provided the ratio of the capillary Bingham number to the Bond number is below a critical value, $\mathcal{J}<1$. We solved the thin-film evolution equation numerically to examine the nonlinear evolution of an initially nearly-flat layer. Numerical simulations were conducted in periodic domains, which allowed for investigation of the initial linear growth of the instability, as well as the subsequent development of a steadily-translating collar. We investigated predictions from the small-$B$ model in the case that there is one steadily-translating collar with a fixed volume, and compared the results to numerical solutions of the full thin-film evolution equation. We found generally good agreement between the structure of the asymptotic model in its various regions with the shapes of $h$ and $Y$ in a numerical simulation at $B=0.02$ (figure 6). We also found generally good agreement between leading-order asymptotic predictions for the collar speed and the corresponding quantity from numerical simulations for a range of $\mathcal{J}$, even when $B$ was not extremely small (figure 5). Including higher-order corrections in the global force balance, arising from contributions from the intermediate regions, resulted in improved predictions for the collar speed. We also used the asymptotic model to show that, when the film thickness ahead of the collar is fixed, increasing $\mathcal{J}$ leads to an increase in both the volume and speed of a steadily-translating collar (figure 4).
@@
Extending our model to describe motion of thick collars or liquid plugs, as has been done for the Newtonian case [4], is left for future studies, as is validation of our results against computational fluid dynamics (CFD) simulations or laboratory experiments with viscoplastic fluids. The finite-time blow-up of slender translating collars that can occur when the film ahead of the collar is sufficiently thick (figure 7) suggests that, in these cases, collars do not remain slender and finite-thickness effects must then be taken into account. It is likely that, where the thin-film theory predicts blow-up, either the collar will saturate as a non-slender collar, will form a liquid plug in the tube, or, if it is on the exterior of the tube, may form a large droplet. In future, finite-thickness effects could be investigated by deriving and analysing a long-wave model, of the type used to predict plug formation in the Rayleigh-Plateau instability previously [15, 16, 27, 32, 25, 46, 28], or by solving the full two-dimensional equations via CFD, which has also been used to study plug formation for Newtonian and non-Newtonian fluids [29, 31, 17, 18] and to study steady viscoplastic plug propagation driven by imposed pressure gradients [12].
There has been some experimental work characterising viscoplastic effects on plug propagation and rupture [47], against which a theoretical model of plug propagation could be compared. It has been shown that plug formation and propagation can induce significant stresses on the tube wall, which may have physiological importance since airway wall epithelial cells can be injured if large enough stresses are exerted on them [19]. In our asymptotic model of quasi-steadily translating collars, a stress approximately equal to the liquid’s yield stress is exerted on the tube wall across the length of the collar. This suggests that taking into account the liquid’s yield stress may be important when considering wall stresses exerted by collar or plug propagation. However, an extended model that can model plug propagation and non-slender collars would be required to confirm this.
We have focused on modelling the formation and motion of individual liquid collars. In numerical simulations, we have used periodic domains with length equal to the most unstable wavelength from linear stability analysis, so that only a single collar can develop in the domain. In longer domains, multiple collars may develop and potentially interact; we leave investigation of the dynamics of interacting viscoplastic collars for future studies.

### v3

@@ IV.3 Intermediate regions
[…]
@@
[…]
@@ IV.4 Bretherton regions
$h=B\mathcal{H}_{\pm}(\xi)+\dots,\quad Y=B\mathcal{H}_{\pm}(\xi)+\dots,$ (61a,b)
where the subscripts again distinguish quantities in the rear and front Bretherton regions. The evolution equation (30) again says that the leading-order flux is constant, with $q=-B^{5/2}\mathcal{U}H_{\infty}^{+}$ in the rear Bretherton region, and $q=-B^{5/2}\mathcal{U}H_{\infty}^{-}$ in the front. Therefore, (30) and ((61a,b)) give
@@
We introduce rescaled variables, defining $X_{\pm}=(3|\mathcal{U}|)^{1/3}\xi/{H_{\infty}^{\pm}}$, and let $\mathcal{H}_{\pm}=H_{\infty}^{\pm}H_{\pm}(X_{\pm})$, so that (62) becomes
@@
$h\sim B^{1/2}H_{2}\hat{X}^{2}\frac{\left(3|\mathcal{U}|\right)^{2/3}}{2H_{\infty}^{+}}+BH_{\infty}^{+}H_{0}+B^{5/4}\frac{2(H_{\infty}^{+})^{2}}{3H_{2}^{2}(3|\mathcal{U}|)^{1/3}\hat{X}}$$\dots$ (66)
@@
$h\sim B^{1/2}\frac{V}{4\pi}\hat{X}^{2}+B\hat{A}_{+}-B^{5/4}\frac{8\pi^{2}\mathcal{U}}{3V^{2}\hat{X}}$$\dots$. (67)
@@
$h\sim B^{1/2}\frac{V}{4\pi}\hat{X}^{2}+B\left(\hat{A}_{+}-2\pi+\frac{4\pi^{2}\mathcal{J}}{V}\right)-B^{5/4}\frac{8\pi^{2}\mathcal{U}}{3V^{2}\hat{X}}$$\dots$. (70)
@@
We have now derived a leading-order expression for $\mathcal{U}$, and relations between $H_{\infty}^{\pm}$, $\mathcal{U}$, $\mathcal{J}$ and $V$. There are further details of the asymptotic model, discussed in appendices B and C, which are required to expose the full structure of the solution. In the front Bretherton region, the layer height, $\mathcal{H}$, has an undulating shape, and there is a sequence of points within the Bretherton region around which $\mathcal{H}^{\prime\prime\prime}\approx 0$. An asymptotically short inner region, in which $Y$ is no longer equal to $h$ at leading order, must exist around each point where $\mathcal{H}^{\prime\prime\prime}=0$. We discuss the structure of these inner regions in appendix B. At the outer edges of both Bretherton regions, $Y$ must approach zero in order for the solution to match onto the uniform unyielded film away from the collar. In appendix C, we discuss the structure of two ‘rigidification regions’ near to the points $\zeta=\zeta_{0}$ and $\zeta=\zeta_{1}$, where $Y\rightarrow 0$. We find that in the rigidification regions, $\zeta-\zeta_{1}=O(B^{1/2})$ and $\zeta_{0}-\zeta=O(B^{1/2})$, respectively, but with $\zeta-\zeta_{1}\ll B^{1/2}\log(1/B)$ and $\zeta_{0}-\zeta\ll B^{1/2}\log(1/B)$, and that the Bretherton region solutions outlined above strictly hold when $\zeta-\zeta_{1}\gg B^{1/2}\log(1/B)$ and $\zeta_{0}-\zeta\gg B^{1/2}\log(1/B)$, respectively. This motivates the $O(B^{1/2}\log(1/B))$ length scales identified in figure 3(b,c) for the distance between the Bretherton regions and the points $\zeta=\zeta_{0}$ and $\zeta=\zeta_{1}$. The rigidification regions resemble analogous regions close to yield surfaces in other thin-film capillary flows [39, 38], but in the current problem, the rigidification regions are always matched onto almost fully-yielded Bretherton regions. Appendix C outlines how the rigidification regions connect the Bretherton regions to the unyielded uniform films away from the collar. None of these details impact the leading-order predictions for $\mathcal{U}$ and $H_{\infty}^{+}$ determined above.
@@ V Translating collars at small $B$
The asymptotic model outlined in §IV provides information on how a single, isolated collar propagates into a uniform film ahead of it when $B\ll 1$ and $\mathcal{J}=O(1)$. Numerical solutions of (22), such as the one shown in figure 2, can also describe the motion of a translating collar when they are run to sufficiently long times to allow for a collar to develop and then begin steadily propagating. Since we enforce periodicity at the sides of our computational domain when solving (22) numerically, the film thicknesses at the front edge and rear edge of a translating collar must be the same. This means that the numerical simulations always reach a state at late times in which the collar is steadily translating. Therefore, we can directly compare predictions from the asymptotic model, in the case that $V$ and $\mathcal{U}$ are constant and $H_{\infty}^{+}=H_{\infty}^{-}$, against late-time predictions from numerical simulations: we do so in §V.1. In §V.2, we discuss how an isolated collar evolves in an unsteady manner when the film thickness ahead and behind are unequal, which can be described by the asymptotic model.
@@ V.1 Steady translation
Figure 4: Dependence of collar volume and speed on Bingham number, $\mathcal{J}$, for fixed values of $H_{\infty}^{-}$. Variation of (a) $V$, and (b) $\mathcal{U}$, for $0<\mathcal{J}<1$, according to (77) and (78). Each solid line corresponds to one value of $H_{\infty}^{-}$, with the values shown being $H_{\infty}^{-}=\{0.4,0.7,1,1.3,1.6\}$. The large arrows indicate increasing values of $H_{\infty}^{-}$. The dashed line in (a) is $V=2\pi\mathcal{J}$, which corresponds to the minimum volume required for the collar to translate.
@@
Figure 5: Data from numerical solutions of (22) with domain length $\mathcal{L}=2\pi/k_{\mathrm{m}}$, for various values of the Bond number, $B$, and Bingham number, $\mathcal{J}=J/B$. (a) Late-time collar speed, $\mathcal{U}=B^{-3/2}U$, evaluated as the gradient of $z_{\max}(t)$ at the end of the simulation, where $z_{\max}$ is the location of the maximum in $h$. (b) $B^{-1}h(z_{\max}\pm\mathcal{L}/2,t_{\mathrm{end}})$, which approximates the deposited film thickness, $H_{\infty}^{-}=B^{-1}h_{\infty}^{-}$, where $t_{\mathrm{end}}$ is the simulation end time. Solid lines show leading-order asymptotic predictions (78) for $\mathcal{U}$ and (77) for $H_{\infty}^{-}$, evaluated by setting $V=\mathcal{L}$. The dashed line in (a) is $\mathcal{U}$ evaluated using (D.18), the global force balance that includes $O(B^{5/4}\log{B})$ corrections to the wall shear stress, for $B=0.025$. The dotted line shows $\mathcal{U}$ determined via (80) and (81), the global force balance in the case that $J=O(B^{3/2})$, again for $B=0.025$. For each $B$, $t_{\mathrm{end}}$ is chosen to be large enough to achieve adequate convergence, ranging from $t_{\mathrm{end}}=8\times 10^{3}$ for $B=0.1$ to $t_{\mathrm{end}}=4\times 10^{4}$ for $B=0.0125$. The initial conditions (9) are used for simulations with $\mathcal{J}=0$. For each $B$, $\mathcal{J}$ is then increased in increments of $0.1$ and the final solution from the previous value of $\mathcal{J}$ is used as the initial conditions; this enables faster convergence to the late-time solutions.
@@
Rather than treating $H_{\infty}^{-}$ as fixed, as we do in figure 4, we could instead treat $V$ as fixed, considering a collar with fixed volume. In this case, from (77) and (78), both the film thickness, $H_{\infty}^{-}$, and the propagation speed, $|\mathcal{U}|$, decrease as $\mathcal{J}$ is increased. The case of fixed $V$ is more amenable to direct comparison with numerical solutions of (22) in periodic domains, since the total volume of fluid within the periodic domain is fixed to be $\mathcal{L}$. The asymptotic prediction for $\mathcal{U}$ compares well with late-time collar speeds in numerical simulations (figure 5a, solid line), as does the prediction for $H_{\infty}^{-}$ when $B$ is sufficiently small (figure 5b). In figure 5(a), there is good (but not perfect) agreement between $B^{-3/2}U$ from numerical simulations, and $\mathcal{U}$, which provides evidence to support the scalings and solutions in the asymptotic theory for small $B$. Even when $B$ is not extremely small, there is reasonably good quantitative agreement between the numerical and asymptotic predictions for translation speed (figure 5), although the asymptotics generally predict a modestly faster speed. At smaller values of $\mathcal{J}$, the speeds from numerical solutions appear to converge towards the asymptotic prediction as $B$ is decreased, but at larger $\mathcal{J}$ there is not clear convergence towards the leading-order asymptotic prediction for the values of $B$ tested in figure 5. When $\mathcal{J}$ is large enough that viscoplastic effects are significant, we may expect that convergence to the small-$B$ asymptotic regime only occurs at much smaller $B$ than when $\mathcal{J}\approx 0$. This is because the largest error term in $Y$ in the main body of the collar is only expected to be $O(B^{1/4})$ larger than the leading-order term ((39a,b)), so $B$ must be very small before the error in $Y$, and by extension the error in the viscous drag in the collar, is negligible.
Figure 5(b) shows a comparison between $H_{\infty}^{-}$ from (77) and $B^{-1}h(z_{\mathrm{max}}\pm\mathcal{L}/2,t_{\mathrm{end}})$, which is an approximation to the deposited film thickness at the end time of numerical simulations. There is a more prominent discrepancy in the deposited film thickness when $B$ is not small for larger values of $\mathcal{J}$, but the agreement generally improves as $B$ is decreased. This suggests that higher-order terms in the small-$B$ expansion for $H_{\infty}^{+}$ may increase in magnitude as $\mathcal{J}$ is increased, meaning the error between leading-order asymptotics and numerics is larger at larger $\mathcal{J}$. Nonetheless, when $B=0.0125$, there is good agreement between numerics and the leading-order approximation for $H_{\infty}^{+}$ for all values of $\mathcal{J}$ tested in figure 5.
For numerical simulations with $\mathcal{J}>1$, there would not be unstable growth and collar formation if the initial conditions (29) were used. However, there is still a subcritical instability that can be triggered if a sufficiently large perturbation is applied to the free surface initially. The asymptotic solution exists for $0\leq\mathcal{J}<V/2\pi$, and steadily-translating collar states are achieved at late-times in simulations throughout this range (figure 5). For $\mathcal{J}>V/2\pi$, we find that there can still be unstable growth in numerical simulations if a large perturbation is applied to the free surface initially (data not shown), but that collars generally appear to decelerate at late times, with the collar speed approaching zero, rather than reaching a steadily-translating state. In the limit $\mathcal{J}\rightarrow(V/2\pi)^{-}$, we have $\mathcal{U}\rightarrow 0^{-}$, so the assumption that the collar speed scales like $U\sim B^{-3/2}$ could break down when $\mathcal{J}$ is sufficiently close to $V/2\pi$. An alternative scaling may be required to fully explore this limit, and a modified asymptotic solution could exist. However, this limit represents only extremely slow collar motion, so we do not explore this in detail, and continue to focus on the results where $\mathcal{U}=O(1)$.
@@
where the $O(B^{5/4}\log(1/B))$ correction arises from contributions from the wall shear stress in the intermediate regions. Replacing the left-hand side of (76) with the expression (79) yields a prediction for the collar speed which generally agrees more closely with numerical results than the leading-order prediction for $\mathcal{U}$ (figure 5a, dashed line).
@@
is a good approximation for $0<\Gamma<20$. Numerical evaluation of $\tilde{F}(\Gamma)$ suggests that $\tilde{F}\approx 2\pi$ when $\Gamma$ is sufficiently large, suggesting that the result (76) for $J=O(B)$ can be recovered in the limit $\hat{J}\rightarrow\infty$. The collar speed determined via (80), with the approximation (81), yields an improved estimate of the steady collar speed when $\mathcal{J}\ll 1$ (figure 5a, dotted line), compared to the leading-order prediction (76) for $\mathcal{U}$. For $B=0.025$, $\Gamma<20$ holds for $0<\mathcal{J}\leq 0.4$, so use of (81) is justified at least for $0<\mathcal{J}\leq 0.4$ in figure 5(a).
@@ VI Discussion
We solved the thin-film evolution equation numerically to examine the nonlinear evolution of an initially nearly-flat layer. Numerical simulations were conducted in periodic domains, which allowed for investigation of the initial growth of the instability, as well as the subsequent development of a steadily-translating collar. We investigated predictions from the small-$B$ model in the case that there is one steadily-translating collar with a fixed volume, and compared the results to numerical solutions of the full thin-film evolution equation. We found generally good agreement between the structure of the asymptotic model in its various regions with the shapes of $h$ and $Y$ in a numerical simulation at $B=0.02$ (figure 6). We also found generally good agreement between leading-order asymptotic predictions for the collar speed and the corresponding quantity from numerical simulations for a range of $\mathcal{J}$, even when $B$ was not extremely small (figure 5a). Including higher-order corrections in the global force balance, arising from contributions from the intermediate regions, resulted in improved predictions for the collar speed. We also used the asymptotic model to show that, when the film thickness ahead of the collar is fixed, increasing $\mathcal{J}$ leads to an increase in both the volume and speed of a steadily-translating collar (figure 4).
@@
Extending our model to describe motion of thick collars or liquid plugs, as has been done for the Newtonian case [5], is left for future studies. Computational fluid dynamics (CFD) simulations have been used to study the problem of pressure-driven liquid plug motion [8], which exhibits some features similar to those of gravity-driven collar motion, including an undulating free surface at the leading edge of the plug, with the shear-rate and shear stress rapidly changing signs where these undulations occur. However, direct comparison of our theoretical results with CFD is left for future studies in which gravity-driven collar motion is simulated. The finite-time blow-up of slender translating collars that can occur when the film ahead of the collar is sufficiently thick (figure 7) suggests that, in these cases, collars do not remain slender and finite-thickness effects must then be taken into account. It is likely that, where the thin-film theory predicts blow-up, either the collar will saturate as a non-slender collar, will form a liquid plug in the tube, or, if it is on the exterior of the tube, may form a large droplet. In future, finite-thickness effects could be investigated by deriving and analysing a long-wave model, of the type used to predict plug formation in the Rayleigh-Plateau instability previously [11, 12, 27, 32, 25, 49, 28], or by solving the full two-dimensional equations via CFD, which has also been used to study plug formation for Newtonian and non-Newtonian fluids. [29, 31, 13, 14]
There has been some experimental work characterising viscoplastic effects on plug propagation and rupture [50], against which a theoretical model of plug propagation could be compared. Bahrani et al. [50] also developed a model of elastoviscoplastic plug propagation, which took into account the thickening of the liquid film deposited behind the plug due to elastic effects. Employing a more complex constitutive law in our model may modify its predictions for the collar dynamics, but this is beyond the scope of the current study. It has been shown that plug formation and propagation can induce significant stresses on the tube wall, which may have physiological importance since airway wall epithelial cells can be injured if large enough stresses are exerted on them [15]. In our asymptotic model of quasi-steadily translating collars, a stress approximately equal to the liquid’s yield stress is exerted on the tube wall across the length of the collar. This suggests that taking into account the liquid’s yield stress may be important when considering wall stresses exerted by collar or plug propagation. However, an extended model that can model plug propagation and non-slender collars would be required to confirm this.
We have focused on modelling the formation and motion of individual liquid collars. In numerical simulations, we have used periodic domains with a fixed length that allows for only a single collar to develop in the domain. In longer domains, multiple collars may develop and potentially interact; we leave investigation of the dynamics of interacting viscoplastic collars for future studies.